OpenStudy (anonymous):

Sarah wants to buy the car that has a better mileage rating. Cars of Company A travel 20 miles per gallon and cars of Company B travel 14 km per liter. [1 mile = 1.6 km; 1 gallon = 3.79 liters] Part A: Which company's car should Sarah buy? Justify your answer by comparing the mileages after converting to the same units. Part B: If 10 gallons of gas is filled in the car which gives better mileage, and you went on a drive and checked on the gas used every 30 miles, describe appropriate scales on the x-axis and y-axis to graph the rate at which the car uses gas.

OpenStudy (anonymous):

\[\frac{ 20~miles }{ 1~gallon }=\frac{ 20*1.6~km }{ 3.79~liters }=\frac{ 20*1.6 }{ 3.79 } km~per~liter \] now compare